# Comprehensive Review and Empirical Evaluation of Causal Discovery Algorithms for Numerical Data

### v1

@@ 4.6 Hybrid Method
One illustrative hybrid approach is the Autoregressive Moving Average - Linear Non-Gaussian Acyclic Model (ARMA-LiNGAM) (Kawahara et al., 2011), which combines Granger Causality (GC) and Structural Equation Modeling (SEM). This composite method resolves both instantaneous and delayed causality by leveraging the attributes of LiNGAM and ARMA models. ARMA-LiNGAM’s integration allows for a more comprehensive analysis of time series data, accommodating both immediate and lagged effects. Incorporating deep learning techniques with GC models, Neural Granger Causality (NGC) (Tank et al., 2021b; Wang et al., 2023) stands out. NGC utilizes the Causal Multilayer Perceptron (CMLP) model to train data, thereby enhancing the accuracy of causal inference tasks. By combining the predictive power of neural networks with the interpretability of Granger causality, NGC offers a robust framework for identifying causal relationships in complex datasets.
@@
Raghu et al. (2018) proposed the Scalable Causation Discovery Algorithm (SCDA), which combines structural equation modeling-based and conditional independence-based methods. SCDA provides a solution for mixed data containing both continuous and discrete sequences. By integrating these two methodologies, SCDA can handle a broader range of data types and improve the robustness of causal inference in heterogeneous datasets.
In summary, hybrid methods in causal discovery leverage the strengths of multiple algorithms to address their respective weaknesses. By combining techniques such as Granger causality, structural equation modeling, neural networks, and information entropy, these hybrid approaches offer powerful tools for uncovering causal relationships in diverse and complex datasets.
@@ 5 Empirical Study Design
The experimental framework is structured across four distinct phases. The inaugural phase involves conducting a comparative assessment of algorithms applied to synthesised datasets with specific features while concurrently evaluating a range of performance metrics. Subsequently, the second phase encompasses presenting and analysing outcomes derived from the initial stage to extract meaningful insights. Transitioning to the third phase, real-world datasets are engaged for testing, utilising the insights garnered in the preceding phase to ascertain the optimal algorithm. This stage aims to verify whether the test results are consistent with the predicted optimal algorithm. The fourth and final phase entails deploying diverse data processing and testing methodologies to ascertain the metadata of the time series datasets. This, in turn, facilitates the extrapolation of insights from the second phase to previously unexplored datasets.
@@ 5.1 Datasets
Given the categorization of testing algorithms into three classes, a concomitant preparation of diverse composite datasets becomes imperative. The first category entails a data generator tasked with producing causal pairs, whereby the data generation tool within the causal discovery toolbox (Kalainathan and Goudet, 2019b) was employed. The second and third categories encompass instantaneous and time-delay causality, for which the data generator in Tigramite framework, posited by Runge et al. (2023a), was harnessed for data synthesis. In the fourth category, characterized by i.i.d. data, we adopted the data generation model established within the gCastle package.
@@ 5.2 Evaluation Metrics
5.3 Algorithms
By collecting resources on GitHub, we introduce in this section the algorithms selected for the experiment, as well as their source code and packages. We mainly use four packages, Causal Discovery Toolbox (CDT) (2019a), gCastle (2021b), causal_discovery_for_time_series (CD_TS)(2022), causal-learn (2024) to implement the testing algorithms.
@@ 5.3 Algorithms
As for the structural equation modeling-based approach, our emphasis lies on exploring a variant of the LiNGAM algorithm (Ikeuchi et al., 2023) tailored for i.i.d. data and time series data, referred to as ICALiNGAM, DirectLiNGAM, RCD, CAM-UV and VARLiNGAM. The TiMINO algorithm is excluded from consideration due to its documented inferiority compared to PCMCI and TCDF (Nauta et al., 2019a).
@@
Structural equation modeling-based ICALiNGAM* LiNGAM; gCastle; causal-learn
@@ 5.4 Environment Settings
5.5 Research Questions
With a focus on the analysis of algorithm performance and result effectiveness, our experiment addresses the following research questions:
RQ 1 is the baseline question in the experiment and a prerequisite for answering other questions. By evaluating the impact of data features on algorithm performance, we establish a foundation that informs subsequent steps in our experimental analysis.
RQ 2 is crucial for determining the effectiveness of insights gained from RQ 1, as it connects the findings from synthetic data to real-world scenarios. This question ensures that our conclusions are not limited to controlled experimental conditions but are also valid in practical applications.
RQ 3 (Generalization to Unknown Datasets):
RQ 3 is a challenging aspect of the experiment as it involves extending the results of RQ 1 and RQ 2 to a broader range of applications. RQ 3.1 focuses on the feasibility of applying our findings to new datasets by analyzing their metadata, ensuring that our methods are robust and versatile. RQ 3.2 aims to translate these validated approaches into practical, user-friendly guidelines that assist practitioners in choosing the best algorithms for their unique datasets, thus bridging the gap between theoretical research and practical implementation.
6 Results Analysis
@@ 6.3 Answer to RQ3: Generalization to Unknown Datasets
To answer RQ 3.1, Table 11 shows that these three metadata can be extracted precisely, with accuracies over $75\%$ and standard deviations of no more than 0.1, indicating that the judgment program is also relatively stable.
@@ 7 Threats to Validity
7.0.1 Threats to internal validity
(1)Correctness of The Codes. A potential threat to the study revolves around the accuracy of the employed codebase. Despite rigorous testing and validation procedures, the complexity of algorithmic implementations and potential oversights during the coding process may give rise to errors. Variations in the code may inadvertently influence the outcomes, thereby introducing a threat to the study’s internal validity.
@@ 7.0.1 Threats to internal validity
(2) Implementation of Control Variables. In light of the comparative analytical methodology employed within the experimental framework of this project, the precision in governing variables assumes paramount importance due to its direct influence on the ensuing output. If multiple testing algorithms are administered across disparate datasets, the potential for consequential impact stemming from inherent dataset dissimilarities becomes salient. Such variations could undermine the cogency of the ultimate conclusions drawn from the study.
@@
7.0.2 Threats to external validity
(1) Metadata selection bias. The external validity of this study could be threatened by metadata selection bias. In light of our methodology, which exclusively constructs testing data predicated upon established causal linkages, noise distribution characteristics, and dataset dimensions, it is prudent to acknowledge the potential bias that crucial time series’ metadata might not be exhaustively contained. Consequently, there remains a prospect for more influential data attributes that could threaten the comprehensiveness and external validity of the experimental outcomes.
@@ 7.0.2 Threats to external validity
(2) Temporal Validity. Considering the incessant evolution characterizing algorithms, it is plausible that our research might not encompass forthcoming optimizations, thereby engendering temporal constraints on the experiment’s conclusions. This scenario presents an external threat to the generalizability of research endeavours.
@@ 8.1 Conclusions and Summary
8.2 Recommendations for Further Work
Opportunities for enhancement and progress persist in this study. Based on the discussion section, a trajectory for future investigations can be charted across three principal dimensions. The initial facet involves considering a broader range of data types, such as high-dimensional datasets or time series incorporating latent variables.
@@ 8.2 Recommendations for Further Work
Additionally, future work could explore the application of causal discovery methods to verify the results of data augmentation (Gao and Li, 2024). For example, these methods could be used to ascertain whether newly generated data retains the causality of the original data. This verification process would enhance the reliability and validity of augmented datasets, ensuring that the fundamental causal relationships are preserved.

### v2

@@ 4.6 Hybrid Method
One illustrative hybrid approach is the Autoregressive Moving Average - Linear Non-Gaussian Acyclic Model (ARMA-LiNGAM) (Kawahara et al., 2011), which combines Granger Causality (GC) and Structural Equation Models (SEM). This composite method resolves both instantaneous and delayed causality by leveraging the attributes of LiNGAM and ARMA models. ARMA-LiNGAM’s integration allows for a more comprehensive analysis of time series data, accommodating both immediate and lagged effects. Incorporating deep learning techniques with GC models, Neural Granger Causality (NGC) (Tank et al., 2021b; Wang et al., 2023) stands out. NGC utilizes the Causal Multilayer Perceptron (CMLP) model to train data, thereby enhancing the accuracy of causal inference tasks. By combining the predictive power of neural networks with the interpretability of Granger causality, NGC offers a robust framework for identifying causal relationships in complex datasets.
@@
Raghu et al. (2018) proposed the Scalable Causation Discovery Algorithm (SCDA), which combines structural equation model-based and conditional independence-based methods. SCDA provides a solution for mixed data containing both continuous and discrete sequences. By integrating these two methodologies, SCDA can handle a broader range of data types and improve the robustness of causal inference in heterogeneous datasets.
Monti et al. (2020) proposed an algorithm called Non-linear SEM Estimation using Non-Stationarity (NonSENS) for bivariate data. This approach employs a deep learning-based method, Time Contrastive Learning (TCL), within a SEM framework, allowing for arbitrary instantaneous nonlinear relationships without assuming additive noise. Notably, the direction of effect in arbitrary nonlinear SEMs is proved to be identifiable (Hyvärinen et al., 2024).
Bystrova et al. (2024) developed two novel algorithms, NBCB and CBNB, which integrate SEM with a constraint-based approach to infer causal graphs from time series. Both approaches are capable of inferring various types of causal graphs including instantaneous and lagged relationships. These algorithms exhibit effectiveness and robustness across both synthetic and real-world datasets.
In summary, hybrid methods in causal discovery leverage the strengths of multiple algorithms to address their respective weaknesses. By combining techniques such as Granger causality, structural equation models, neural networks, and information entropy, these hybrid approaches offer powerful tools for uncovering causal relationships in diverse and complex datasets.
@@ 5 Empirical Study Design
The aim of this section is to design experiments to answer the three RQs in section 1. RQ1 is comparison of algorithm performance, RQ 2 is real-world applicability, and RQ3 is generalization to unknown datasets.
In light of this, the experimental framework is structured across four distinct phases. The inaugural phase involves conducting a comparative assessment of algorithms applied to synthesised datasets with specific features while concurrently evaluating a range of performance metrics. Subsequently, the second phase encompasses presenting and analysing outcomes derived from the initial stage to extract meaningful insights. Transitioning to the third phase, real-world datasets are engaged for testing, utilising the insights garnered in the preceding phase to ascertain the optimal algorithm. This stage aims to verify whether the test results are consistent with the predicted optimal algorithm. The fourth and final phase entails deploying diverse data processing and testing methodologies to ascertain the metadata of the time series datasets. This, in turn, facilitates the extrapolation of insights from the second phase to previously unexplored datasets.
@@ 5.1 Datasets
Given the categorization of testing algorithms into three classes, a concomitant preparation of diverse composite datasets becomes imperative. The first category entails a data generator tasked with producing causal pairs, whereby the data generation tool within the causal discovery toolbox (Kalainathan et al., 2020) was employed. The second and third categories encompass instantaneous and time-delay causality, for which the data generator in Tigramite framework, posited by Runge et al. (2023a), was harnessed for data synthesis. In the fourth category, characterized by i.i.d. data, we adopted the data generation model established within the gCastle package (Zhang et al., 2021b).
@@ 5.2 Evaluation Metrics
5.3 Experimental Algorithms
By collecting resources on GitHub, we introduce in this section the algorithms selected for the experiment, as well as their source code and packages. We mainly use four packages, Causal Discovery Toolbox (CDT) (2019), gCastle (2021a), Causal Discovery for Time Series (CD_TS)(2022), causal-learn (2024a) to implement the testing algorithms.
@@ 5.3 Experimental Algorithms
As for the structural equation model-based approach, our emphasis lies on exploring a variant of the LiNGAM algorithm (Ikeuchi et al., 2023) tailored for i.i.d. data and time series data, referred to as ICALiNGAM, DirectLiNGAM, RCD, and VARLiNGAM. The TiMINO algorithm is excluded from consideration due to its documented inferiority compared to PCMCI and TCDF (Nauta et al., 2019a).
@@
Structural equation model-based ICALiNGAM* LiNGAM; gCastle; causal-learn
@@ 5.4 Environment Settings
6 Expirical Results and Analyses
@@ 6.3 Answer to RQ3: Generalization to Unknown Datasets
To answer RQ 3.1, Table 13 shows that these three metadata can be extracted precisely, with accuracies over 75% and standard deviations of no more than 0.1, indicating that the judgment program is also relatively stable.
@@ 7 Threats to Validity
7.1 Threats to Internal Validity
7.1.1 Correctness of The Codes
A potential threat to the study revolves around the accuracy of the employed codebase. Despite rigorous testing and validation procedures, the complexity of algorithmic implementations and potential oversights during the coding process may give rise to errors. Variations in the code may inadvertently influence the outcomes, thereby introducing a threat to the study’s internal validity.
@@ 7.1.1 Correctness of The Codes
7.1.2 Implementation of Control Variables
In light of the comparative analytical methodology employed within the experimental framework of this project, the precision in governing variables assumes paramount importance due to its direct influence on the ensuing output. If multiple testing algorithms are administered across disparate datasets, the potential for consequential impact stemming from inherent dataset dissimilarities becomes salient. Such variations could undermine the cogency of the ultimate conclusions drawn from the study.
@@ 7.1.2 Implementation of Control Variables
7.2 Threats to external validity
7.2.1 Metadata selection bias
The external validity of this study could be threatened by metadata selection bias. In light of our methodology, which exclusively constructs testing data predicated upon established causal linkages, noise distribution characteristics, and dataset dimensions, it is prudent to acknowledge the potential bias that crucial time series’ metadata might not be exhaustively contained. Consequently, there remains a prospect for more influential data attributes that could threaten the comprehensiveness and external validity of the experimental outcomes.
@@ 7.2.1 Metadata selection bias
7.2.2 Temporal Validity
Considering the incessant evolution characterizing algorithms, it is plausible that our research might not encompass forthcoming optimizations, thereby engendering temporal constraints on the experiment’s conclusions. This scenario presents an external threat to the generalizability of research endeavours.
@@ 8.1 Conclusions
8.2 Discussion on Future Research Directions
Opportunities for enhancement and progress persist in this study. Based on the discussion section, a trajectory for future investigations can be charted across three principal dimensions to improve the algorithm guideline system. The initial facet involves considering a broader range of data types, such as high-dimensional datasets or time series incorporating latent variables.
@@ 8.2 Discussion on Future Research Directions
Besides, it is important to discuss research directions related to existing research gaps in the field, mainly focusing on three aspects:
  • Datasets and algorithms: One significant issue is the limited availability of real datasets with ground truth, which restricts algorithm experiments to only a few real datasets. Therefore, developing a broader range of real datasets is crucial for benchmarking causal discovery algorihthms. Additionally, in the development of new algorithms, two-step approaches show significant potential. For instance, integrating causal discovery methods with neural networks in different stages could overcome the limitations relying on a single method.
  • Algorithm selector: A critical research direction is the joint analysis of data features and causal graph structures. This effort aims to develop a comprehensive algorithm selector (Runge et al., 2023a) capable of addressing all possible data scenarios, thereby assisting users in handling various complex situations.
  • Causality or Association? Understanding the fundamental differences between causality and regression is crucial. Conducting empirical research to clarify the quantitative differences between causal algorithms and association-based methods will provide valuable insights for advancing causal research and statistical techniques. This understanding is also essential for evaluating the impact that causal learning could bring to the field of explainable AI (Montavon et al., 2018).
Additionally, future work could explore the application of causal discovery methods to verify the results of data augmentation (Gao et al., 2023). For example, these methods could be used to ascertain whether newly generated data retains the causality of the original data. This verification process would enhance the reliability and validity of augmented datasets, ensuring that the fundamental causal relationships are preserved.